A hollow metal sphere is made of metal 4mm thick and has external diameter of 12cm. If the density of the metal is 8.8g/cm cube,what is the mass of the sphere
Answer by mananth(16946)


Volume of metal is External volume - Internal volume
= 4/3 * pi* r1^3-4/3 * pi r2 ^3
= 4/3 * pi * (r1^3-r2^3)
= 4/3 *3.14 * (6^3-5.6^3)
volume of metal = 169 cm^3
density = 8.8 gm/cm^3
weight = 169.07 * 8.8 =1488 grams